# Universal constant order fluctuations for the cokernels of block lower triangular matrices

### v1

@@ 1 Introduction
where $\operatorname{RowSpace}(\mathbf{M})$ is the subgroup of $\mathbb{Z}^{n}$ generated by the rows of $\mathbf{M}$. When we study the cokernels of random matrices, we are often interested in the Sylow $p$-subgroup of the cokernel. There is a widespread universality phenomenon for the cokernels of random matrices, that is, in many cases the limiting behavior of the Sylow $p$-subgroup of the cokernel does not depend on the finer details of the given random matrix model. The simplest form of this universality was proved by Wood [26]: Given $\varepsilon>0$ and a prime $p$, we say that a $\mathbb{Z}$-valued random variable $X$ is $(p,\varepsilon)$-balanced if $\mathbb{P}(X\equiv r\mod p)\leq 1-\varepsilon$ for all $r\in\mathbb{Z}$. Wood proved that if $\mathbf{M}_{n}$ is a random $n\times n$ matrix with independent $(p,\varepsilon)$-balanced entries, then the distribution of the Sylow $p$-group of $\operatorname{cok}(\mathbf{M}_{n})$ converges to the Cohen-Lenstra distribution. The Cohen-Lenstra distribution is a distribution on the finite abelian $p$-groups, where the probability assigned to a group $G$ is proportional to $|\operatorname{Aut}(G)|^{-1}$. This distribution first appeared in conjectures about the distribution of class groups of quadratic number fields [7]. Note that this limiting distribution does not depend on the specific distribution of the entries. Nguyen and Wood [21] weakened the assumptions of the theorem above: It is enough to assume that the entries of $\mathbf{M}_{n}$ are $(p,\varepsilon_{n})$-balanced where $\varepsilon_{n}$ converges to 00 sufficiently slowly. The proofs of these theorems rely on the moment method of Wood for random finite abelian groups [25, 27].
@@
Nguyen and Van Peski [19, 18] considered matrix products of the form $\mathbf{A}_{1}\mathbf{A}_{2}\cdots\mathbf{A}_{k}$, where $\mathbf{A}_{1},\dots,\mathbf{A}_{k}$ are i.i.d. $n\times n$ random matrices, where the entries are i.i.d. copies of some $\mathbb{Z}$-valued random variable which is non-constant mod $p$. Let $\Gamma$ be the Sylow $p$-group of $\operatorname{cok}(\mathbf{A}_{1}\mathbf{A}_{2}\cdots\mathbf{A}_{k})$. For fixed $k$, Nguyen and Van Peski showed that as $n\to\infty$, $\Gamma$ has a limiting distribution [19]. This limiting distribution is a modified version of the Cohen-Lenstra distribution where the probability assigned to a group $G$ is obtained by also taking into account the number of chains of subgroups of $G$ of length $k$. When $k$ goes to infinity together with $n$, the size of $\Gamma$ goes to infinity. Thus, $\Gamma$ can not have a limiting distribution in the usual sense. However, Nguyen and Van Peski [18] proved that if $n,k\to\infty$ such that $k=O(e^{(\log n)^{1-\varepsilon}})$, then the fluctuation of $\Gamma$ stays of constant order. They also determined the limiting distribution of these fluctuations. The limit again does not depend on the distribution of the entries. See Theorem 3 for a more detailed statement. This is a new interesting phenomenon, so it is a natural question whether we can see the same type of constant order fluctuations for other random groups of growing order.
@@
$\mathbf{C}=\begin{pmatrix}\mathbf{A}_{1}&\mathbf{0}&\ddots&&&\\ \mathbf{I}&\mathbf{A}_{2}&\mathbf{0}&\ddots&\\ \mathbf{0}&\mathbf{I}&\mathbf{A}_{3}&\mathbf{0}&\ddots&\\ \ddots&\mathbf{0}&\mathbf{I}&\mathbf{A}_{4}&\mathbf{0}&\ddots&\\ &\ddots&\ddots&\ddots&\ddots&\ddots\\ &&\ddots&\mathbf{0}&\mathbf{I}&\mathbf{A}_{k}\end{pmatrix}.$
@@
see Section A.1. Thus, $\operatorname{cok}(\mathbf{C})$ shows the same behavior as the cokernel of matrix products. Given the prevalence of the universality phenomenon in the study of cokernels of random matrices, one could expect the same behavior from many other random block lower triangular matrix models. Theorem 2 below provides such examples of block lower triangular matrices. Therefore, the hidden block lower triangular structure of matrix products gives a more general context where the constant order fluctuations found by Nguyen and Van Peski can be observed universally.
@@
Combining Theorem 1 with a special case of the rescaled moment method of Nguyen and Van Peski [18] stated in Theorem 8, one can describe the limiting fluctuations of the Sylow $p$-subgroup of $\operatorname{cok}(\mathbf{C}^{(h)})$, as we explain next.
@@
Thus, the Sylow $p$-subgroup of the cokernel of an integral random matrix, which is a finite abelian $p$-group can be also represented by a partition, or equivalently by the corresponding Young diagram. Thus, a random integral matrix determines a random Young diagram. Theorem 2 below describes the joint fluctuations of the first $d$ columns of the Young diagram corresponding the Sylow $p$-group of $\operatorname{cok}(\mathbf{C}^{(h)})$. Before we can state this theorem, we need a few further notions.
@@
$\operatorname{Sig}_{d}$ =$=\{(\lambda_{1},\lambda_{2},\dots,\lambda_{d})\in\mathbb{Z}^{d}\,:\,\lambda_{1}\geq\lambda_{2}\cdots\geq\lambda_{d}\}\text{ and }$
$\operatorname{Sig}_{d}^{\geq 0}$ =$=\{(\lambda_{1},\lambda_{2},\dots,\lambda_{d})\in\mathbb{Z}^{d}\,:\,\lambda_{1}\geq\lambda_{2}\cdots\geq\lambda_{d}\geq 0\}.$
@@
${\mathbb{E}p^{\left\langle\mathcal{L}_{d,p^{-1},\chi}\,,\,\lambda\right\rangle}}=\frac{((p-1)\chi)^{|\lambda|}}{|\lambda|!}c(G_{\lambda^{\prime}},|\lambda|)\qquad\text{ for all }\lambda\in\operatorname{Sig}_{d}^{\geq 0}.$
@@
Comparing Theorem 2 with the theorem of Nguyen and Van Peski [18] below, we see that the cokernel of block lower triangular matrices lies in the same universality class as the cokernel of matrix products.
Theorem 3 (Nguyen and Van Peski [18]).
@@
Finally, we provide an improvement on Theorem 3 by replacing the condition on $k_{n}$ with a weaker one.
Theorem 4.
Theorem 3 remains valid, if we replace the condition $k_{n}=O(e^{(\log n)^{1-\varepsilon}})$ with
@@ 2 Preliminaries
Theorem 2 follows from Theorem 1 by using the following theorem of Nguyen and Van Peski [18].
Theorem 8.
@@
3 The proof of Theorem 1
@@
If $\langle\mathbf{g}_{i-1}\rangle\not\subseteq\langle\mathbf{g}_{i}\rangle$, then by Lemma 7, we have the estimate
@@
Also by (A.3), we have
$\mathbb{P}\left(\mathbf{A}_{i,i}\mathbf{g}_{i}=-\mathbf{A}_{i,i-1}\mathbf{g}_{i-1}-\sum_{j=1}^{i-2}\mathbf{B}^{\prime}_{i,j}\mathbf{g}_{j}\,\Big{|}\,\mathbf{A}_{i,i-1}\mathbf{g}_{i-1}+\sum_{j=1}^{i-2}\mathbf{B}^{\prime}_{i,j}\mathbf{g}_{j}\in\langle\mathbf{g}_{i}\rangle^{n_{i}}\right)\leq\max_{\mathbf{f}\in\langle\mathbf{g}_{i}\rangle^{n_{i}}}\mathbb{P}(\mathbf{A}_{i,i}\mathbf{g}_{i}=\mathbf{f}).$
Thus, with the notation
$w=\left|\left\{i\in\{2,3,\dots,k\}\,:\,\langle\mathbf{g}_{i-1}\rangle\not\subseteq\langle\mathbf{g}_{i}\rangle\right\}\right|,$
we have
$\mathbb{P}(\mathbf{C}\mathbf{g}=0\,|\,\mathbf{B}=\mathbf{B}^{\prime})\leq(1-\varepsilon)^{wn_{\bullet}}\prod_{i=1}^{n}\max_{\mathbf{f}\in\langle\mathbf{g}_{i}\rangle^{n_{i}}}\mathbb{P}(\mathbf{A}_{i,i}\mathbf{g}_{i}=\mathbf{f}).$
@@
$\mathbb{P}(\mathbf{C}\mathbf{g}=0)\leq(1-\varepsilon)^{wn_{\bullet}}\prod_{i=1}^{n}\max_{\mathbf{f}\in\langle\mathbf{g}_{i}\rangle^{n_{i}}}\mathbb{P}(\mathbf{A}_{i,i}\mathbf{g}_{i}=\mathbf{f}).$ (6)
@@
$W(\mathbf{H})=\left\{i\in\{2,3,\dots,k\}\,:\,{H}_{i-1}\not\subseteq{H}_{i}\right\}\text{ and }w(\mathbf{H})=|W(\mathbf{H})|.$
@@
$\leq 2(1-\varepsilon)^{w(\mathbf{H})n_{\bullet}}.$ by (A.7) (7)
@@
$T(\mathbf{H})=\left\{i\in\{1,2,\dots,k\}\,:\,{H}_{i-1}\subsetneq{H}_{i}\right\}\text{ and }t(\mathbf{H})=|T(\mathbf{H})|.$
@@
Thus, one can recover $\mathbf{H}$ from $W(\mathbf{H})\cup T(\mathbf{H})$ and $\mathbf{H}\restriction(W(\mathbf{H})\cup T(\mathbf{H}))$. Therefore, it follows that for any $u$, we have
@@
Combining (8) and Claim 9, we see that for any $w\geq 1$, we have
@@
$\sum_{\begin{subarray}{c}\mathbf{g}\in G^{n}\\ w(\operatorname{Type}(\mathbf{g}))=w\end{subarray}}\mathbb{P}(\mathbf{C}\mathbf{g}=0)\leq 6\frac{(|\operatorname{Sub}(G)|k)^{2(\ell(G)+1)w}}{w!}(1-\varepsilon)^{wn_{\bullet}},$
@@
$\sum_{\begin{subarray}{c}\mathbf{g}\in G^{n}\\ w(\operatorname{Type}(\mathbf{g}))\geq 1\end{subarray}}\mathbb{P}(\mathbf{C}\mathbf{g}=0)$ $\leq 6\sum_{w=1}^{\infty}\frac{\left((|\operatorname{Sub}(G)|k)^{2(\ell(G)+1)}(1-\varepsilon)^{n_{\bullet}}\right)^{w}}{w!}$
@@
$\lim_{h\to\infty}\sum_{\begin{subarray}{c}\mathbf{g}\in G^{n(h)}\\ w(\operatorname{Type}(\mathbf{g}))\geq 1\end{subarray}}\mathbb{P}(\mathbf{C}^{(h)}\mathbf{g}=0)=0.$ (9)
@@
Note that if for a $\mathbf{g}\in G^{n}$, we have $w(\operatorname{Type}(\mathbf{g}))=0$, then for all $i\geq 2$, we have
@@
We already seen that
@@
$\sum_{\begin{subarray}{c}\mathbf{g}\in G^{n}\\ w(\operatorname{Type}(\mathbf{g}))=0\\ t(\operatorname{Type}(\mathbf{g}))=i\end{subarray}}\mathbb{P}(\mathbf{C}\mathbf{g}=0)=(1+o_{h}(1))c(G,i){{k}\choose{i}}=(1+o_{h}(1))c(G,i)\frac{k^{i}}{i!}.$
@@
$\lim_{h\to\infty}(k(h))^{-\ell(G)}\sum_{\begin{subarray}{c}\mathbf{g}\in G^{n(h)}\\ w(\operatorname{Type}(\mathbf{g}))=0\\ t(\operatorname{Type}(\mathbf{g}))=i\end{subarray}}\mathbb{P}(\mathbf{C}^{(h)}\mathbf{g}=0)=\begin{cases}\frac{c(G,\ell(G))}{\ell(G)!}&\text{if }i=\ell(G),\\ 0&\text{otherwise}.\end{cases}$ (11)
Using Lemma 5, we see that
@@
$=k^{-\ell(G)}\left(\sum_{\begin{subarray}{c}\mathbf{g}\in G^{n}\\ w(\operatorname{Type}(\mathbf{g}))\geq 1\end{subarray}}\mathbb{P}(\mathbf{C}\mathbf{g}=0)+\sum_{i=0}^{\ell(G)}\sum_{\begin{subarray}{c}\mathbf{g}\in G^{n}\\ w(\operatorname{Type}(\mathbf{g}))=0\\ t(\operatorname{Type}(\mathbf{g}))=i\end{subarray}}\mathbb{P}(\mathbf{C}\mathbf{g}=0)\right).$
Combining this with (9) and (11), Theorem 1 follows.
4 The proof of Theorem 4
The proof of Theorem 4 is very similar to the proof of Theorem 1. It is actually even easier, so we only sketch the details. Let $G$ be a finite abelian $p$-group. We drop the index of $k_{n}$. Let $\mathbf{A}_{1},\mathbf{A}_{2},\dots,\mathbf{A}_{k}$ be independent ($(\{p\},\varepsilon)$-balanced $n\times n$ random matrices. We have
@@ 4 The proof of Theorem 4
$\sum_{\begin{subarray}{c}\mathbf{g}_{1}\in H_{1}^{n}\\ \langle\mathbf{g}_{1}\rangle=H_{1}\end{subarray}}\sum_{\begin{subarray}{c}\mathbf{g}_{2}\in H_{2}^{n}\\ \langle\mathbf{g}_{2}\rangle=H_{2}\end{subarray}}\cdots\sum_{\begin{subarray}{c}\mathbf{g}_{k}\in H_{k}^{n}\\ \langle\mathbf{g}_{k}\rangle=H_{k}\end{subarray}}\prod_{i=1}^{k}\mathbb{P}(\mathbf{A}_{i}\mathbf{g}_{i}=\mathbf{g}_{i-1})$ $\leq\sum_{\begin{subarray}{c}\mathbf{g}_{1}\in H_{1}^{n}\\ \langle\mathbf{g}_{1}\rangle=H_{1}\end{subarray}}\sum_{\begin{subarray}{c}\mathbf{g}_{2}\in H_{2}^{n}\\ \langle\mathbf{g}_{2}\rangle=H_{2}\end{subarray}}\cdots\sum_{\begin{subarray}{c}\mathbf{g}_{k}\in H_{k}^{n}\\ \langle\mathbf{g}_{k}\rangle=H_{k}\end{subarray}}\prod_{i=1}^{k}\max_{\mathbf{f}\in H_{i}^{n}}\mathbb{P}(\mathbf{A}_{i}\mathbf{g}_{i}=\mathbf{f})$
@@
where the last inequality follows from Lemma 6. Similarly,
@@
The following along the lines of the proof of Theorem 1, it can be proved that
@@
Thus, Theorem 4 follows by using Theorem 8.

### v2

# Universal constant order fluctuations for the cokernels of block triangular matrices

@@ 1 Introduction
where $\operatorname{RowSpace}(\mathbf{M})$ is the subgroup of $\mathbb{Z}^{n}$ generated by the rows of $\mathbf{M}$. When we study the cokernels of random matrices, we are often interested in the Sylow $p$-subgroup of the cokernel. There is a widespread universality phenomenon for the cokernels of random matrices, that is, in many cases the limiting behavior of the Sylow $p$-subgroup of the cokernel does not depend on the finer details of the given random matrix model. The simplest form of this universality was proved by Wood [26]: Given $\varepsilon>0$ and a prime $p$, we say that a $\mathbb{Z}$-valued random variable $X$ is $(p,\varepsilon)$-balanced if $\mathbb{P}(X\equiv r\mod p)\leq 1-\varepsilon$ for all $r\in\mathbb{Z}$. Wood proved that if $\mathbf{M}_{n}$ is a random $n\times n$ matrix with independent $(p,\varepsilon)$-balanced entries, then the distribution of the Sylow $p$-group of $\operatorname{cok}(\mathbf{M}_{n})$ converges to the Cohen-Lenstra distribution. The Cohen-Lenstra distribution is a distribution on the finite abelian $p$-groups, where the probability assigned to a group $G$ is proportional to $|\operatorname{Aut}(G)|^{-1}$. This distribution first appeared in conjectures about the distribution of class groups of quadratic number fields [7]. Note that this limiting distribution does not depend on the specific distribution of the entries. Nguyen and Wood [21] weakened the assumptions of the theorem above: It is enough to assume that the entries of $\mathbf{M}_{n}$ are $(p,\varepsilon_{n})$-balanced where $\varepsilon_{n}$ converges to 0 sufficiently slowly. The proofs of these theorems rely on the moment method of Wood for random finite abelian groups [25, 27].
@@
Nguyen and Van Peski [19, 18] considered matrix products of the form $\mathbf{A}_{1}\mathbf{A}_{2}\cdots\mathbf{A}_{k}$, where $\mathbf{A}_{1},\dots,\mathbf{A}_{k}$ are i.i.d. $n\times n$ random matrices, where the entries are i.i.d. copies of some $\mathbb{Z}$-valued random variable which is non-constant mod $p$. Let $\Gamma$ be the Sylow $p$-group of $\operatorname{cok}(\mathbf{A}_{1}\mathbf{A}_{2}\cdots\mathbf{A}_{k})$. For fixed $k$, Nguyen and Van Peski showed that as $n\to\infty$, $\Gamma$ has a limiting distribution [19]. This limiting distribution is a modified version of the Cohen-Lenstra distribution where the probability assigned to a group $G$ is obtained by also taking into account the number of chains of subgroups of $G$ of length $k$. When $k$ goes to infinity together with $n$, the size of $\Gamma$ goes to infinity. Thus, $\Gamma$ can not have a limiting distribution in the usual sense. However, Nguyen and Van Peski [18] proved that if $n,k\to\infty$ such that $k=O(e^{(\log n)^{1-\varepsilon}})$, then the fluctuation of $\Gamma$ stays of constant order. They also determined the limiting distribution of these fluctuations. The limit again does not depend on the distribution of the entries. See Theorem 1.3 for a more detailed statement. This is a new interesting phenomenon, so it is a natural question whether we can see the same type of constant order fluctuations for other random groups of growing order.
@@
$\mathbf{C}=\begin{pmatrix}\mathbf{A}_{1}&\mathbf{0}&\ddots&&&\\ \mathbf{I}&\mathbf{A}_{2}&\mathbf{0}&\ddots&\\ \mathbf{0}&\mathbf{I}&\mathbf{A}_{3}&\mathbf{0}&\ddots&\\ \ddots&\mathbf{0}&\mathbf{I}&\mathbf{A}_{4}&\mathbf{0}&\ddots&\\ &\ddots&\ddots&\ddots&\ddots&\ddots\\ &&\ddots&\mathbf{0}&\mathbf{I}&\mathbf{A}_{k}\end{pmatrix}$. (1)
@@
see Section A.1. Thus, $\operatorname{cok}(\mathbf{C})$ shows the same behavior as the cokernel of matrix products. Given the prevalence of the universality phenomenon in the study of cokernels of random matrices, one could expect the same behavior from many other random block lower triangular matrix models. Theorem 1.2 below provides such examples of block lower triangular matrices. Therefore, the hidden block lower triangular structure of matrix products gives a more general context where the constant order fluctuations found by Nguyen and Van Peski can be observed universally.
@@
Combining Theorem 1.1 with a special case of the rescaled moment method of Nguyen and Van Peski [18] stated in Theorem 2.4, one can describe the limiting fluctuations of the Sylow $p$-subgroup of $\operatorname{cok}(\mathbf{C}^{(h)})$, as we explain next.
@@
Thus, the Sylow $p$-subgroup of the cokernel of an integral random matrix, which is a finite abelian $p$-group can be also represented by a partition, or equivalently by the corresponding Young diagram. Thus, a random integral matrix determines a random Young diagram. Theorem 1.2 below describes the joint fluctuations of the first $d$ columns of the Young diagram corresponding the Sylow $p$-group of $\operatorname{cok}(\mathbf{C}^{(h)})$. Before we can state this theorem, we need a few further notions.
@@
$\operatorname{Sig}_{d}$ $=\{(\lambda_{1},\lambda_{2},\dots,\lambda_{d})\in\mathbb{Z}^{d}\,:\,\lambda_{1}\geq\lambda_{2}\cdots\geq\lambda_{d}\}\text{ and }$
$\operatorname{Sig}_{d}^{\geq 0}$ $=\{(\lambda_{1},\lambda_{2},\dots,\lambda_{d})\in\mathbb{Z}^{d}\,:\,\lambda_{1}\geq\lambda_{2}\cdots\geq\lambda_{d}\geq 0\}.$
@@
For two vectors $v,w\in\mathbb{Z}^{d}$, their inner product is defined as $\langle v,w\rangle=\sum_{i=1}^{d}v_{i}w_{i}$.
@@
${\mathbb{E}p^{\left\langle\mathcal{L}_{d,p^{-1},\chi}\,,\,\lambda\right\rangle}}=\frac{((p-1)\chi)^{|\lambda|}}{|\lambda|!}c(G_{\lambda^{\prime}},|\lambda|)\hskip 23.49976pt\text{ for all }\lambda\in\operatorname{Sig}_{d}^{\geq 0}.$
@@
Comparing Theorem 1.2 with the theorem of Nguyen and Van Peski [18] below, we see that the cokernel of block lower triangular matrices lies in the same universality class as the cokernel of matrix products.
Theorem 1.3 (Nguyen and Van Peski [18]).
@@
Although Theorem 1.2 does not cover the matrices obtained from matrix products by the construction given in (1), with only small modifications, our methods can also be applied to such matrices. In fact, our proof provides an improvement on Theorem 1.3 by replacing the condition on $k_{n}$ with a weaker one.
Theorem 1.4.
Theorem 1.3 remains valid, if we replace the condition $k_{n}=O(e^{(\log n)^{1-\varepsilon}})$ with
@@
Acknowledgments: The author is grateful to the anonymous referees for their useful suggestions. The author was supported by the KKP 139502 project and Dynasnet European Research Council Synergy project – grant number ERC-2018-SYG 810115.
@@ 2 Preliminaries
Theorem 1.2 follows from Theorem 1.1 by using the following theorem of Nguyen and Van Peski [18].
Theorem 2.4.
@@
3 The proof of Theorem 1.1
3.1 The outline of the proof
@@
We define
$w(\mathbf{g})$ $=\left|\left\{i\in\{2,3,\dots,k\}\,:\,\langle\mathbf{g}_{i-1}\rangle\not\subseteq\langle\mathbf{g}_{i}\rangle\right\}\right|,$
$t(\mathbf{g})$ $=\left|\left\{i\in\{1,2,\dots,k\}\,:\,\langle\mathbf{g}_{i-1}\rangle\subsetneq\langle\mathbf{g}_{i}\rangle\right\}\right|,$
with the convention that $\langle\mathbf{g}_{0}\rangle=\{0\}$.
By Lemma 2.1, we need to find the asymptotics of the sum
$\sum_{\mathbf{g}\in G^{n}}\mathbb{P}(\mathbf{C}\mathbf{g}=0).$ (5)
We split the sum according to the values of $w(\mathbf{g})$ and $t(\mathbf{g})$, and prove the following statements:
  • The vectors $\mathbf{g}\in G^{n}$ such that $w(\mathbf{g})>0$ have a negligible contribution to the sum in (5). See equation (11).
  • If $w(\mathbf{g})=0$, then $t(\mathbf{g})$ can take the values $0,1,\dots,\ell(G)$. Given an $i\in\{0,1,\dots,\ell(G)\}$, the total contribution of the vectors $\mathbf{g}\in G^{n}$ such that $w(\mathbf{g})=0$ and $t(\mathbf{g})=i$ to the sum is
    $(1+o(1))c(G,i)\frac{k^{i}}{i!},$
    see equation (13). Here, the choice of $i=\ell(G)$ gives us the dominant contribution.
Once we prove these statements, Theorem 1.1 follows easily.
3.2 Details of the proof
@@
If $\langle\mathbf{g}_{i-1}\rangle\not\subseteq\langle\mathbf{g}_{i}\rangle$, then by Lemma 2.3, we have the estimate
@@
By (A.3), the matrix $\mathbf{A}_{i,i}$ is independent from $\mathbf{A}_{i,i-1}$. Therefore,
$\mathbb{P}\left(\mathbf{A}_{i,i}\mathbf{g}_{i}=-\mathbf{A}_{i,i-1}\mathbf{g}_{i-1}-\sum_{j=1}^{i-2}\mathbf{B}^{\prime}_{i,j}\mathbf{g}_{j}\,\Big|\,\mathbf{A}_{i,i-1}\mathbf{g}_{i-1}+\sum_{j=1}^{i-2}\mathbf{B}^{\prime}_{i,j}\mathbf{g}_{j}\in\langle\mathbf{g}_{i}\rangle^{n_{i}}\right)\\ \leq\max_{\mathbf{f}\in\langle\mathbf{g}_{i}\rangle^{n_{i}}}\mathbb{P}(\mathbf{A}_{i,i}\mathbf{g}_{i}=\mathbf{f}).$
Thus,
$\mathbb{P}(\mathbf{C}\mathbf{g}=0\,|\,\mathbf{B}=\mathbf{B}^{\prime})\leq(1-\varepsilon)^{w(\mathbf{g})n_{\bullet}}\prod_{i=1}^{n}\max_{\mathbf{f}\in\langle\mathbf{g}_{i}\rangle^{n_{i}}}\mathbb{P}(\mathbf{A}_{i,i}\mathbf{g}_{i}=\mathbf{f}).$
@@
$\mathbb{P}(\mathbf{C}\mathbf{g}=0)\leq(1-\varepsilon)^{w(\mathbf{g})n_{\bullet}}\prod_{i=1}^{n}\max_{\mathbf{f}\in\langle\mathbf{g}_{i}\rangle^{n_{i}}}\mathbb{P}(\mathbf{A}_{i,i}\mathbf{g}_{i}=\mathbf{f}).$ (8)
@@
[…]
Note that the definition $w(\mathbf{H})$ is consistent with our earlier definition of $w(\mathbf{g})$ in the sense that $w(\mathbf{g})=w(\operatorname{Type}(\mathbf{g}))$.
@@
$\sum_{\mathbf{g}\in\operatorname{Type}^{-1}(\mathbf{H})}$ $\mathbb{P}(C\mathbf{g}=0)$
@@
$\leq 2(1-\varepsilon)^{w(\mathbf{H})n_{\bullet}}.$ (by (A.7)) (9)
@@
[…]
@@
Thus, one can recover $\mathbf{H}$ from $W(\mathbf{H})\cup T(\mathbf{H})$ and the restriction of $\mathbf{H}$ to $W(\mathbf{H})\cup T(\mathbf{H})$. Therefore, it follows that for any $u$, we have
@@
Combining (10) and Claim 3.1, we see that for any $w\geq 1$, we have
@@
$\sum_{\begin{subarray}{c}\mathbf{g}\in G^{n}\\ w(\mathbf{g})=w\end{subarray}}\mathbb{P}(\mathbf{C}\mathbf{g}=0)\leq 6\frac{(|\operatorname{Sub}(G)|k)^{2(\ell(G)+1)w}}{w!}(1-\varepsilon)^{wn_{\bullet}},$
@@
$\sum_{\begin{subarray}{c}\mathbf{g}\in G^{n}\\ w(\mathbf{g})\geq 1\end{subarray}}\mathbb{P}(\mathbf{C}\mathbf{g}=0)$ $\leq 6\sum_{w=1}^{\infty}\frac{\left((|\operatorname{Sub}(G)|k)^{2(\ell(G)+1)}(1-\varepsilon)^{n_{\bullet}}\right)^{w}}{w!}$
@@
$\lim_{h\to\infty}\sum_{\begin{subarray}{c}\mathbf{g}\in G^{n(h)}\\ w(\mathbf{g})\geq 1\end{subarray}}\mathbb{P}(\mathbf{C}^{(h)}\mathbf{g}=0)=0.$ (11)
@@
Note that if for a $\mathbf{g}\in G^{n}$, we have $w(\mathbf{g})=0$, then for all $i\geq 2$, we have
@@
We have already seen that
@@
$\sum_{\begin{subarray}{c}\mathbf{g}\in G^{n}\\ w(\mathbf{g})=0\\ t(\mathbf{g})=i\end{subarray}}\mathbb{P}(\mathbf{C}\mathbf{g}=0)=(1+o_{h}(1))c(G,i){{k}\choose{i}}=(1+o_{h}(1))c(G,i)\frac{k^{i}}{i!}.$ (13)
@@
$\lim_{h\to\infty}(k(h))^{-\ell(G)}\sum_{\begin{subarray}{c}\mathbf{g}\in G^{n(h)}\\ w(\mathbf{g})=0\\ t(\mathbf{g})=i\end{subarray}}\mathbb{P}(\mathbf{C}^{(h)}\mathbf{g}=0)=\begin{cases}\frac{c(G,\ell(G))}{\ell(G)!}&\text{if }i=\ell(G),\\ 0&\text{otherwise}.\end{cases}$ (14)
Using Lemma 2.1, we see that
@@
$=k^{-\ell(G)}\left(\sum_{\begin{subarray}{c}\mathbf{g}\in G^{n}\\ w(\mathbf{g})\geq 1\end{subarray}}\mathbb{P}(\mathbf{C}\mathbf{g}=0)+\sum_{i=0}^{\ell(G)}\sum_{\begin{subarray}{c}\mathbf{g}\in G^{n}\\ w(\mathbf{g})=0\\ t(\mathbf{g})=i\end{subarray}}\mathbb{P}(\mathbf{C}\mathbf{g}=0)\right).$
Combining this with (11) and (14), Theorem 1.1 follows.
4 The proof of Theorem 1.4
The proof of Theorem 1.4 is very similar to the proof of Theorem 1.1. It is actually even easier, so we only sketch the details. Let $G$ be a finite abelian $p$-group. We drop the index of $k_{n}$. Let $\mathbf{A}_{1},\mathbf{A}_{2},\dots,\mathbf{A}_{k}$ be independent $(\{p\},\varepsilon)$-balanced $n\times n$ random matrices. We have
@@
$\sum_{\begin{subarray}{c}\mathbf{g}_{1}\in H_{1}^{n}\\ \langle\mathbf{g}_{1}\rangle=H_{1}\end{subarray}}\sum_{\begin{subarray}{c}\mathbf{g}_{2}\in H_{2}^{n}\\ \langle\mathbf{g}_{2}\rangle=H_{2}\end{subarray}}\cdots\sum_{\begin{subarray}{c}\mathbf{g}_{k}\in H_{k}^{n}\\ \langle\mathbf{g}_{k}\rangle=H_{k}\end{subarray}}$ $\prod_{i=1}^{k}\mathbb{P}(\mathbf{A}_{i}\mathbf{g}_{i}=\mathbf{g}_{i-1})$
$\leq\sum_{\begin{subarray}{c}\mathbf{g}_{1}\in H_{1}^{n}\\ \langle\mathbf{g}_{1}\rangle=H_{1}\end{subarray}}\sum_{\begin{subarray}{c}\mathbf{g}_{2}\in H_{2}^{n}\\ \langle\mathbf{g}_{2}\rangle=H_{2}\end{subarray}}\cdots\sum_{\begin{subarray}{c}\mathbf{g}_{k}\in H_{k}^{n}\\ \langle\mathbf{g}_{k}\rangle=H_{k}\end{subarray}}\prod_{i=1}^{k}\max_{\mathbf{f}\in H_{i}^{n}}\mathbb{P}(\mathbf{A}_{i}\mathbf{g}_{i}=\mathbf{f})$
@@
where the last inequality follows from Lemma 2.2. Similarly,
@@
Then following along the lines of the proof of Theorem 1.1, it can be proved that
@@
Thus, Theorem 1.4 follows by using Theorem 2.4.